OpenStudy (anonymous):

(cos3x)^2+(sin3x)^2=cos(f(x)) find f(x)

OpenStudy (anonymous):

\[ (\cos3x)^2+(\sin3x)^2=1 \]

OpenStudy (anonymous):

So I guess...\[ 1=\cos(f(x)) \\ \arccos(1) = f(x) \]

OpenStudy (anonymous):

It's a very strange problem.